# Polish Academy of Sciences

### v2

@@ 5.4 Slab-based Filtering
The four-dimensional coordinates are then projected into 3D space using a standard perspective projection. Avatars are generated as three-dimensional objects based on visual features stored in the $\mathbf{X}_{\text{visual}}$ matrix and rendered using the tools provided by the dpVision platform (Figure 8). To ensure interactive performance with large datasets, the avatars are deliberately simplified in geometry, allowing a large number of them to be rendered and manipulated in real time.
Figure 8: Part of a sample swarm of avatars with Chernoff-like faces perspectively projected into 3D space.
6 TEST DATA
Given the difficulty of perceiving multivariate data, we need a test data set with artificially created n-dimensional points ($n>3$) whose projection onto any combination of 3 or 2 dimensions is known and recognizable on sight.
One well-known 4D solid is the tesseract [12], a four-dimensional hypercube. However, as a regular hypersolid, it’s not the best choice for our demonstrative purposes, as its orthogonal 3D projection or cross-section in any of its constituent coordinate planes would look exactly the same. We chose to disturb this symmetry by changing its aspect ratio from $1:1:1:1$ to $3:4:6:9$, a small-integer approximation of a chain of three golden ratios. It was displayed in dpVision, and a programmed procedure rotated it around five coordinate planes (in four dimensions, rotation occurs around a plane rather than a line).
Figure 9 shows the projection, into the 3D viewing space of dpVision, of this hypersolid in one of its rotated positions.
Figure 9: The anisotropic tesseract rotating in 4D, projected into a 3D viewing space.
7 NAVIGATING THE DATA
The four coordinate 3D spaces of the 4D hyperspace exhibit a topology homeomorphic to that of the tetrahedron. Each of them can be transformed into each of the other ones via a $90^{\circ}$ rotation (in 4D) around one of the six coordinate planes, just like each face of a tetrahedron can be transformed into any other one by a rotation (in 3D) of approximately $71^{\circ}$ around one of the six edges. The principle is shown in Figure 10, limited to four rotations (rather than the full six) to avoid clutter. The video file accompanying this paper shows our test set performing, in steps of $10^{\circ}$ at a time, the following drill:
Figure 10: Rotating the view in 4D.
  1. 1.
    Attribute numbers to coordinate spaces:
    1. (a)
      $XYZ$
    2. (b)
      $XYT$
    3. (c)
      $XZT$
    4. (d)
      $YZT$
  2. 2.
    make first coordinate space the viewing space; place two of its axes parallel to the edges of the screen.
  3. 3.
    in steps of $10^{\circ}$, rotate the cloud in 3D $90^{\circ}$ around the axis parallel to the vertical edges of the screen, then to the horizontal.
  4. 4.
    rotate the cloud in 4D, also in nine $10^{\circ}$ steps, around one of the coordinate planes, so that the next coordinate 3D space coincides with the viewing space
  5. 5.
    repeat from step 3 until all the coordinate spaces have been viewed.
This example is only intended to demonstrate the geometric aspects of our approach. A fuller illustration uses the red wines file of the Vinho Verde wine dataset.
Figure 11: Example of visualization of the data from the Vinho Verde wine dataset.
In this test (Figure 11), the pH and residual sugar content were mapped to the X and Y dimensions, respectively, while dimensions Z and T represented the highest-energy plane across a hyperspace spanned by citric acid, chlorides, total sulfur dioxide, and sulfates. Overall quality of the wine was represented by the smile (or lack thereof) on the Chernoff face, and alcohol content by the length of its nose. The density (specific weight) of the wines was represented by the color of the face, and it can be observed that three distinct densities prevail in the data set, leading to a small set of colors out of a potentially near-continuous scale of hues.
8 CONTRIBUTIONS
This work provides several contributions to the field of multidimensional data visualization:
  1. 1.
    Semantic division of data dimensions
    We propose a novel conceptual framework that divides the dimensions of a dataset into two categories:
    • dimensions that are intuitively understandable to human users (e.g., quality ratings, sensory characteristics),
    • and more abstract or technical dimensions (e.g., chemical or process parameters).
    Unlike previous methods, which typically treat all dimensions equally, we assign intuitive dimensions to visual features of avatars (shape, facial expression, color) and project abstract dimensions into 3D spatial coordinates after possible dimensionality reduction.
  2. 2.
    Integration of avatar-based glyphs with interactive 3D navigation
    While Chernoff faces have historically been used in static 2D representations, we integrate avatar glyphs into a fully interactive 3D environment. Users can freely explore the data cloud, rotate the viewpoint, adjust the visible slab, and examine clusters and patterns dynamically.
  3. 3.
    Practical implementation as an open-source plugin
    We implemented our method as a working plugin for the open-source platform dpVision. The tool supports user-driven mapping of dimensions to visual properties and interactive spatial exploration, offering immediate practical applicability.
  4. 4.
    Demonstration on both synthetic and real-world datasets
    We validated our approach using two datasets:
    • a synthetic 4D anisotropic hypercube (for controlled geometric validation),
    • and the Vinho Verde wine dataset (a 12-dimensional real-world dataset), illustrating the practical benefits of the method in an applied context.
  5. 5.
    Foundations for future extensibility
    The visualization system is designed with modularity and flexibility in mind. Its architecture supports alternative avatar types, adjustable mappings of data dimensions, and dynamic filtering. While this version targets moderately sized datasets, its structure provides a basis for scaling to larger data collections, as further discussed in the Future Work section.
@@ 9 CONCLUSION
In this paper, we presented a novel method for visualizing high-dimensional data by combining two natural strengths of the human visual system: the understanding of 3D space and the recognition of facial traits.
Our approach introduces an important distinction between intuitive and abstract data dimensions and maps them differently for optimal interpretability: intuitive features are assigned to avatar characteristics, while abstract features are reduced and projected into 3D space.
@@
We implemented the method as a plugin for the open-source dpVision platform and demonstrated its effectiveness on synthetic and real-world datasets. Preliminary results suggest that this combination of avatar-based glyphs and interactive 3D navigation offers an intuitive and flexible way to explore complex data structures, particularly when users are not domain experts in mathematics.
Further development directions are outlined in the next section and include formal usability studies, exploration of alternative glyph styles to mitigate emotional bias, automated feature mapping, and scalability improvements to handle larger datasets.
By distinguishing between different types of dimensions and tailoring their visual representation, our plugin proposes a human-centered paradigm that could benefit many fields where complex data must be understood and interpreted visually.
10 FUTURE WORK
There are several directions in which this work can be further developed. One natural extension is to allow for different types of avatars. While Chernoff-like faces take advantage of the human ability to perceive expressions, they also carry emotional connotations that may bias the interpretation. Therefore, alternative glyphs—such as neutral geometric shapes or biologically inspired forms like trees—could be offered as interchangeable avatar styles, depending on the context of the data and the preferences of the user. Such alternatives could reduce the emotional bias associated with human-like faces and allow more neutral data interpretation depending on the application domain.
Another area of development is the automation of dimension assignment. At present, the mapping of data dimensions to either spatial coordinates or avatar features is performed manually. In future versions, this process could be assisted or fully automated using heuristics or machine learning methods that take into account feature importance, variance, or user-defined priorities.
A further line of improvement concerns scalability. While the current implementation is effective for moderately sized datasets (typically up to a few thousand points), its perceptual clarity decreases as the number of avatars grows. Glyph-based methods such as ours are not ideally suited for very large-scale datasets, as visual clutter and cognitive overload quickly become limiting factors. Future versions could integrate mechanisms such as dynamic filtering, clustering, sampling, or zoomable detail views to maintain interpretability in larger data collections. Hybrid solutions that combine glyphs with statistical overviews may also be explored.
Finally, a formal evaluation of the system is planned. Although the visualization method appears promising and has shown good results in exploratory settings, its usability and effectiveness should be validated through controlled user studies in various application domains. These studies could include tasks such as anomaly detection, group comparison, or classification support, with participants from different backgrounds and levels of expertise.
11 SUPPLEMENTARY MATERIAL
The source code for the plugin is available online under the LGPL 2.1 licence at https://github.com/iitis/N-Dim-view. The source code for the dpVision framework is also available online at https://github.com/pojdulos/dpVision.

### v3

@@ 3.2.4 Slab-based Filtering
The four-dimensional coordinates are then projected into 3D space using a standard perspective projection. Avatars are generated as three-dimensional objects based on visual features stored in the $\mathbf{X}_{\text{visual}}$ matrix and rendered using the tools provided by the dpVision platform. To ensure interactive performance with large datasets, the avatars are deliberately simplified in terms of geometry, allowing a large number of them to be rendered and manipulated in real time.
4 Results
In this section, we present experimental results on both synthetic and real-world datasets to illustrate the interpretability and practical utility of the proposed avatar-based multidimensional visualization method. The analyses focus on (i) how intuitive and technical variables are assigned and visualized, (ii) what insights spatial avatars provide compared to classical approaches, and (iii) the impact of variable assignment choices.
4.1 Small-scale illustration: Synthetic politicians dataset (N=12)
This illustrative example utilizes a manually constructed synthetic dataset of 12 fictional politicians, each described by nine features that reflect both voter perception and performance metrics. Intuitive features (number of promises, fulfillment rate, public sympathy, general popularity) were mapped to avatar visual traits (e.g., smile, nose length, skin color). In contrast, technical features (sympathy, economic views, social views) were assigned to spatial coordinates (X, Y, Z axes). The remaining variables (media activity, voting effectiveness, age) were left anonymous for PCA. The variable group_numeric encodes four archetypes (populists, conservatives, liberals, technocrats), used only for color coding of avatar hair; each sample is labeled by name.
Figure 5: Comparison of classical visualization methods for the synthetic politicians dataset. Top left: PCA scatter plot (first two components); Top right: t-SNE projection; Bottom left: parallel coordinates plot; Bottom right: Chernoff faces grid (4$\times$3) for all twelve politicians, colored and labeled by group. Each approach reveals different aspects of the data: scatter plots show grouping structure, parallel coordinates display multivariate profiles, and Chernoff faces summarize intuitive features—but none allows simultaneous interpretation of both structure and profile as directly as spatial avatars (see Fig. 6).
To provide a fair comparison with established visualization techniques, Figure 5 presents the same dataset using four classical methods: PCA scatter plot, t-SNE projection, parallel coordinates, and a 4$\times$3 grid of Chernoff faces. While PCA and t-SNE reveal the grouping structure, and parallel coordinates allow for the comparison of individual profiles, none of these methods offers the same level of simultaneous, intuitive insight into both position and feature profiles as spatial avatars. In particular, Chernoff faces communicate intuitive attributes but lack spatial context; conversely, scatter plots provide structure but obscure the source of differences between cases.
For the twelve synthetic politicians, classical methods highlight some group separation (notably in PCA and t-SNE), but interpreting individual differences remains challenging. The Chernoff faces grid allows for visual comparison of intuitive features, yet it does not convey the geometric relationships between politicians or facilitate the detection of outliers in attribute space. This contrast underscores the unique advantage of spatial avatars in combining the interpretability of individual profiles with structural awareness.
Variable assignment:
Avatar features: number of promises (skin color), promise fulfillment (nose length), general popularity (smile), group affiliation (hair color). Spatial coordinates: sympathy (X), economic views (Y), social views (Z). Anonymous/PCA: media activity, voting effectiveness, age.
Figure 6: Spatial avatar visualization of the synthetic politicians dataset. Avatar features encode intuitive traits (skin color: number of promises; nose length: promise fulfillment; smile: general popularity; hair color: group). Spatial coordinates reflect sympathy (X), economic views (Y), and social views (Z). Names and group colors aid interpretation. This visualization enables direct comparison of political archetypes, identification of outliers, and exploration of individual profiles.
Interpretation and interactive exploration:
The spatial avatar visualization enables immediate and intuitive comparison of both the position and the profile of each fictional politician. Clusters and polarization become visible in space, while characteristic features (such as promise fulfillment or popularity) are instantly ”readable” from the avatar’s face. Group affiliation is recognized by hair color. Importantly, the tool supports interactive exploration, allowing users to rotate, zoom, and navigate the scene to examine clusters, compare individuals, and discover patterns from multiple perspectives. This interactivity facilitates flexible hypothesis testing and nuanced understanding—features not available in static PCA scatter plots.
PCA loadings and spatial interpretation:
Principal Component Analysis was performed under two scenarios: (top) only on anonymous features (not mapped to avatars or spatial coordinates); (bottom) on all features assigned to spatial coordinates or left anonymous (all except avatar features). Figure 7 presents heatmaps of PCA loadings and the corresponding scatter plots.
Figure 7: Effect of variable assignment on principal component analysis for the synthetic politicians dataset. (top) PCA on anonymous features only. (bottom) PCA on all features assigned to spatial coordinates or left anonymous (all except avatar features). Variable abbreviations: MA—media activity, VE—voting effectiveness, AG—age, SY—sympathy, EV—economic views, SV—social views, SP—sympathy.
The PCA scatter plot in Figure 5 (top left) provides a classical projection of the entire feature set for all politicians, serving as a baseline for comparison. In contrast, Figures 7a and b show PCA performed on, respectively, (a) only anonymous features not mapped to avatars or spatial coordinates and (b) all features mapped to spatial coordinates or left anonymous. These alternative PCA projections illustrate how the interpretability of the spatial arrangement depends critically on which features are included in the dimensionality reduction.
In both PCA scenarios, the principal components are determined by mixtures of technical variables (e.g., media activity, voting effectiveness, age). No single feature fully dominates the spatial arrangement; both principal components represent blends of the anonymous and/or spatial features. Crucially, attributes such as promise fulfillment or group affiliation, which are easily interpreted in the avatar-based view, are not visible in the PCA scatter plot geometry unless they are explicitly included in the projection. It highlights the interpretability advantage of spatial avatars: only by explicit mapping can the user guarantee that key variables are visually represented.
Insights and sensitivity analysis:
Changing the assignment of features (e.g., switching "fulfillment rate" from an avatar trait to a spatial coordinate or PCA) alters both the visual arrangement and interpretability of the visualization. Spatial avatars thus enable flexible, user-driven mapping, supporting either interpretability or structural discovery as needed.
4.2 Medium-scale illustration: Synthetic soft drinks dataset (N=100)
The synthetic soft drinks dataset comprises 100 fictional beverages generated to simulate market diversity with three main product groups (classic, premium, and extreme). Each beverage is described by intuitive sensory attributes (sweetness, fizziness, overall rating, and color intensity) and technical features (price and caffeine). The group_numeric variable encodes the product segment, used only for color coding; each sample is also labeled by name.
Figure 8: Comparison of classical visualization methods for the synthetic soft drinks dataset. Top left: PCA scatter plot; Top right: t-SNE projection; Bottom left: parallel coordinates plot; Bottom right: Chernoff faces grid (6$\times$4) for a subset of beverages, colored by group. While PCA and t-SNE reveal some group structure, and parallel coordinates enable visualization of feature distributions, Chernoff faces offer a glimpse into sensory profiles of selected products. However, none of these methods provides a direct, interpretable synthesis of both product similarity and profile, as does the spatial avatar approach.
To provide context for the effectiveness of our approach, Figure 8 shows the synthetic drinks dataset visualized using four classical techniques: PCA scatter plot, t-SNE projection, parallel coordinates, and a Chernoff faces grid for selected products. PCA and t-SNE projections highlight clear clustering by product type, with the principal components and t-SNE axes separating the main beverage groups. However, interpreting which specific features drive the separation or identifying outliers based on concrete sensory traits remains challenging. Parallel coordinates enable a detailed comparison of feature values but quickly become cluttered at this scale. Chernoff faces communicate some intuitive traits yet are difficult to interpret en masse and do not encode structural relationships between products.
Variable assignment:
Intuitive features (sweetness, fizziness, overall rating, color intensity) were mapped to avatar visual traits (e.g., smile, nose length, skin color, hair color). Technical features (price, caffeine) were used for spatial coordinates (X and Y axes). The remaining attributes—sourness, sugar, citric acid, CO2 pressure, preservatives, and color intensity—were included as anonymous variables for PCA-based analysis. This assignment is reflected in the figures below.
Figure 9: Spatial avatar visualization of the synthetic soft drinks dataset. Avatar features encode intuitive sensory attributes: sweetness (skin color), fizziness (nose length), overall rating (smile curvature), and product group (hair color). Spatial coordinates reflect technical properties: price (X) and caffeine (Y). This visualization enables immediate comparison of product types and detection of unusual profiles, with both group structure and individual differences readily interpretable.
PCA loadings and spatial interpretation:
Principal component analysis was performed under two scenarios: (top) only on anonymous features (not mapped to avatars or spatial coordinates), and (bottom) on all features assigned to spatial coordinates or left anonymous. Figure 10 presents heatmaps of PCA loadings and corresponding scatter plots for both cases.
Figure 10: Effect of variable assignment on principal component analysis for the synthetic soft drinks dataset. (top) PCA on anonymous features only. (bottom) PCA on all features assigned to spatial coordinates or left anonymous (all except avatar features). In both cases, intuitive variables such as overall rating are not directly represented in the geometry unless they are included in the feature set, and their effect may only be visible through color. Variable abbreviations: SOU—sourness, SUG—sugar, CIT—citric acid, CO2—CO2 pressure, PRE—preservatives, COL—color intensity, PRC—price, CAF—caffeine.
Notably, in both PCA scenarios, clear clusters emerge corresponding to beverage groups, as principal components are determined by mixtures of compositional variables (e.g., sourness, sugar, citric acid, preservatives). However, while structural groupings are evident, the interpretation of which sensory or technical attributes drive the group separation and direct identification of outliers remains less intuitive than in the spatial avatar visualization (Fig. 9), where both group membership and profile details are visible at a glance.
Insights and sensitivity analysis:
This example demonstrates that, although dimensionality reduction techniques like PCA and t-SNE can recover the underlying group structure when relevant features dominate the variance, only the spatial avatar approach allows for the immediate interpretation of which specific sensory traits differentiate products, providing direct, profile-based insight. Switching a feature, such as ”sweetness” or ”fizziness”, from an avatar trait to an anonymous (PCA) feature alters the PCA scatter plot and principal component structure but at the expense of immediate interpretability in the avatar view. This flexibility allows users to prioritize either structural discovery or interpretability as needed.
4.3 Large-scale real-world: Wine dataset (N=1599)
We evaluated the method on the Vinho Verde dataset [5, 6], which contains 1,599 samples of Portuguese red wines, each characterized by 12 physicochemical and sensory attributes measured in laboratory analysis. This dataset is widely used as a benchmark for multivariate modeling and classification tasks in oenology.
Figure 11: Classical projection methods for the wine dataset. Left: PCA scatter plot colored by quality rating; Right: t-SNE projection colored by quality. For large datasets such as this, other classical methods (e.g., parallel coordinates or glyph-based Chernoff faces) rapidly become uninterpretable due to severe overplotting and visual clutter and are therefore omitted.
Classical visualization techniques such as PCA and t-SNE scatter plots (Figure 11) provide a global view of data structure. However, for the wine dataset, no distinct clusters or quality-based groupings are apparent; samples of differing quality are dispersed throughout both projections. Other approaches, such as parallel coordinates or glyph-based Chernoff faces, are not shown, as they become uninformative for $N\gg 100$ due to overplotting and loss of interpretability.
Variable assignment:
Quality, alcohol content, and residual sugar were assigned as avatar features (e.g., smile, face color, nose length), as these are familiar and interpretable for non-experts. Four technical variables (volatile acidity, citric acid, residual sugar, sulfates) were assigned to spatial coordinates (X, Y, Z, T). All remaining technical features were omitted from the spatial mapping in this scenario.
Figure 12: Spatial avatar visualization of the Vinho Verde wine dataset. Avatar features encode quality, alcohol content, and residual sugar; spatial coordinates reflect volatile acidity, citric acid, residual sugar, and sulfates. This mapping enables immediate identification of wines with unusual or extreme property profiles, as well as direct interpretation of both spatial position and visual traits.
PCA loadings and spatial interpretation:
For comparison, principal component analysis was performed on all technical features except those assigned to avatar traits. Figure 13 presents a heatmap of PCA loadings and the corresponding scatter plot, with samples colored by quality.
Figure 13: Principal component analysis for the wine dataset based on all technical variables except avatar-assigned ones. The first two principal components are mixtures of several physicochemical properties (PC1: volatile acidity 0.57, density 0.40, sulfates 0.44; PC2: residual sugar 0.59, citric acid 0.52, density 0.45). No clear clusters or quality-based groupings emerge—high- and low-quality wines are scattered across the projection. It highlights the limitation of PCA-based projections for interpretability compared to explicit, user-driven assignments in spatial avatars. Variable abbreviations: VAC—volatile acidity, CAC—citric acid, CLO—chlorides, FSD—free sulfur dioxide, SUL—sulfates, DEN—density, SUG—residual sugar.
In the PCA projection, the principal components reflect combinations of technical features, and the spatial arrangement does not reveal distinct clusters or gradients by quality. By contrast, the spatial avatar visualization (Fig. 12) ensures that critical consumer-facing attributes remain visible and interpretable due to explicit user assignment of axes.
Insights and sensitivity analysis:
Assigning ”quality” or ”alcohol” as a spatial or PCA feature, rather than as an avatar trait, would affect the composition of principal components but reduce the immediate interpretability of these key attributes—demonstrating the practical benefit of flexible, user-driven assignment in the proposed approach.
5 Discussion
This study introduces a flexible and interpretable framework for multidimensional data visualization, placing the user at the center of the exploratory process. The core innovation lies in the semantic division of variables: Users can assign the most meaningful interpretable features (such as quality, popularity, or sensory attributes) to avatar characteristics, while more abstract or technical variables (e.g., chemical or behavioral parameters) are mapped to spatial coordinates. Unlike traditional approaches, such as PCA or parallel coordinates—which treat all variables equally and often obscure critical features—our method allows for user-driven, context-sensitive mapping that supports practical interpretability.
Spatial avatars further extend classical glyph-based visualization by embedding expressive, user-configurable avatars in a fully interactive 3D environment. Users can dynamically explore the data cloud, rotate and zoom into clusters, and rapidly test hypotheses by reassigning features to avatars or axes. This interactive capacity enables both the detection of structural patterns and the immediate comparison of individual profiles—capabilities not readily available in static projection-based visualizations.
Empirical results across synthetic and real-world datasets—including fictional politicians, synthetic soft drinks, and the Vinho Verde wine collection—demonstrate that spatial avatars make group structure, outliers, and archetypes directly observable and interpretable. In contrast, classical PCA-based projections often mix multiple technical variables into principal components, rendering key features invisible or ambiguous from the user’s perspective.
By empowering users to control the mapping of features and explore data interactively, the spatial avatar framework bridges the gap between statistical rigor and practical, human-centered interpretation. The system, implemented as an open-source plugin for the dpVision platform, is adaptable to various analytical goals, data types, and user preferences.
5.1 Limitations
While the proposed approach offers significant advantages, several limitations should be acknowledged. The semantic assignment of variables to ”intuitive” or ”technical” categories relies on user expertise and task context, introducing some subjectivity; different users or domains may lead to different choices and, thus, different interpretations. The number of avatar visual features is limited, restricting the range of intuitive variables that can be displayed directly, which may necessitate dimensionality reduction or the omission of some information. When many variables are treated as ”anonymous” and incorporated via PCA, interpretability can be lost since principal components represent linear combinations rather than original features. Moreover, expressive avatars may introduce emotional or cognitive biases, and for very large or high-dimensional datasets, visual clutter or computational limits could arise. Finally, variable assignment is currently manual; integrating automated or user-guided assignment tools remains an important direction for further work.
Although a formal user study has not yet been conducted, the method has been informally evaluated among collaborating researchers. These preliminary trials indicated that certain avatar features—such as face color or hair color—are readily distinguished and effective for communicating differences, while others, including smile curvature or subtle facial traits, may be less easily perceived. This observation highlights the importance of meticulous glyph design and perceptual testing to enhance interpretability. Future work will involve systematic user studies to quantitatively assess the recognizability of individual avatar features and guide further refinements of the visual encoding.
Addressing these limitations is a natural path for development, including more robust assignment strategies, alternative glyphs to reduce bias, and enhanced support for complex or large-scale datasets.
5.2 Advantages and utility of spatial avatars
Compared to conventional dimensionality reduction methods, the spatial avatar approach provides direct, user-controlled interpretability by allowing the explicit mapping of intuitive features to avatar traits. It supports immediate visual recognition of unusual or exemplary cases and leverages innate human pattern recognition skills. Technical variables projected onto spatial coordinates reveal underlying structures, clusters, or outliers, while avatars simultaneously communicate key user-relevant attributes. Interactive mapping and dynamic exploration enable users to iteratively test hypotheses, adjust variable assignments, and visualize the impact on the data structure.
Empirical case studies (see Section Results) show that spatial avatars reveal patterns and support interpretations that are less accessible or even hidden in standard visualization techniques such as PCA. Sensitivity analyses demonstrate how variable assignment impacts interpretability and cluster formation, underscoring the method’s flexibility and value across diverse analytical contexts.
5.3 Future work
Several directions for further development are envisaged. One approach is to expand the repertoire of avatar types: while Chernoff-like faces take advantage of human abilities to perceive facial traits, alternative glyphs—such as neutral geometric shapes or biologically inspired forms—could reduce emotional bias or adapt to new domains. Another area is the automation of variable assignment; future versions could integrate heuristics or machine learning to suggest optimal mapping based on data properties or user priorities. Scalability remains an open challenge: dynamic filtering, clustering, sampling, or zoomable detail views may help maintain interpretability in larger data collections, possibly in combination with statistical overviews. Ultimately, formal usability studies are necessary to validate the approach in controlled settings and to inform refinements for diverse application areas.
@@ 6 Conclusion
In this paper, we presented a novel method for visualizing high-dimensional data by integrating two natural strengths of human perception: spatial understanding and intuitive recognition of visual patterns. Our approach distinguishes between intuitive and abstract data dimensions, mapping them in different ways for optimal interpretability. Intuitive features are assigned to avatar characteristics, while technical features are reduced and projected into 3D space.
@@
Implemented as a plugin for the open-source dpVision platform, the method was demonstrated on synthetic and real-world datasets. The combination of avatar-based glyphs and interactive 3D navigation provides an intuitive and flexible way to explore complex data structures, facilitating the discovery of patterns and individualized interpretation.
Future work will focus on formal usability evaluation, the development of alternative glyph styles, automated variable mapping, and further improvements to scalability. By enabling users to tailor visual representation to their needs, our framework advances a human-centered paradigm for the visual analysis of complex multidimensional data.
Author Contributions
Conceptualization, L.L.; methodology, L.L.; software, L.L. and D.P.; validation, L.L. and D.P.; formal analysis, L.L.; investigation, D.P.; resources, L.L.; data curation, D.P.; writing—original draft preparation, L.L. and D.P.; writing—review and editing, L.L. and D.P.; visualization, D.P.; supervision, L.L.; project administration, D.P. All authors have read and agreed to the published version of the manuscript.
Funding
This research received no external funding.
Data Availability Statement
The source code for the plugin is available at https://github.com/iitis/N-Dim-view. All synthetic datasets used in this study, together with the additional code required to reproduce the experiments and visualizations, are included in this repository. The real-world wine dataset used is publicly available and formally cited in the References. The source code for the dpVision framework is available at https://github.com/pojdulos/dpVision.
Conflicts of Interest
The authors declare no conflict of interest.
Acknowledgments
The authors would like to thank Krzysztof Domino, PhD, for his valuable advice and motivation, which supported the creation of this article.
List of Abbreviations
PCA Principal Component Analysis
PC Principal Component
t-SNE t-distributed Stochastic Neighbor Embedding
dpVision Open-source Data Visualization Platform
GUI Graphical User Interface
SO(2) Special Orthogonal Group in 2D (if used in text)
SO(4) Special Orthogonal Group in 4D (if used in text)
SVD Singular Value Decomposition (if used in text)